There are four red, three green, and five blue discs. Find the probability that the two discs of same color are drawn out


1
Expert's answer
2021-12-27T04:05:15-0500
"4+3+5=12"


"P(R,R)=\dfrac{4}{12}\cdot\dfrac{3}{11}=\dfrac{1}{11}"

"P(G,G)=\dfrac{3}{12}\cdot\dfrac{2}{11}=\dfrac{1}{22}"

"P(B,B)=\dfrac{5}{12}\cdot\dfrac{4}{11}=\dfrac{5}{33}"

"P(2\ same)=P(R,R)+P(G,G)+P(B, B)"

"=\dfrac{1}{11}+\dfrac{1}{22}+\dfrac{5}{33}=\dfrac{19}{66}"